# Improving entanglement purification through coherent superposition of roles

### v1

@@ III.2 Imperfect operations
The cSWAP operation, also known as the Fredkin gate [57, 58], has been extensively explored in the literature, and various experimental proposals for its high-fidelity implementation already exist [59, 60, 61, 62, 63, 64, 65]. These proposals are based on alternative strategies that circumvent the limitations of decomposing the cSWAP gate into elementary gates. Alternatively, the superposed EPP scheme could be implemented in an interferometric-based way [49, 51], circumventing those problems and possibly enhancing the quality of the implementation of such a gate.
@@
Figs. 4a4b show the operational regime and yields of the protocols under amplitude damping noise, where the noise strength is assumed to be the same for the cNOTs ($q_{cnot}$) and cSWAPs operations ($q_{cSWAP}$). Notably, the superposed strategy, despite the noise from the cSWAP operations, tolerates higher levels of operational noise compared to the double selection protocol, both clearly surpassing basic approaches. Additionally, the superposed EPP outperforms the double selection protocol in terms of fidelity evolution and overall performance.
Fig. 5a shows the operational regimes under local depolarizing noise and initial Werner states, for different values of the cSWAP noisy strength. We remark that the unique features of this gate could lead to alternative improved practical implementations [59, 60, 61, 62, 63, 64, 65, 49, 51].
Finally, we analyze in Fig. 5a the protocols performance in fidelity under special noise considerations, motivated by realistic modeling of current quantum hardware [53]. Following such a model, operational noise is defined as a concatenation of local depolarizing, damping, and dephasing noise affecting each qubit after each gate, i.e., $\xi_{\rm damp}\circ\xi_{\rm damp}\circ\xi_{\rm deph}$. The same noise parameters are used for every gate, consistently showing an advantage of the superposed EPP.
Relevant and consistent improvements are found under different considerations of initial states and noisy operations.
@@
IV Further extensions
The tools presented in this work can offer a recursive method for constructing more efficient EPPs, which can be seamlessly integrated with existing entanglement purification strategies.
We illustrate and analyze various examples herein, emphasizing the advantages derived from our approach. Importantly, our findings suggest that superposed entanglement purification and distillation approaches have the potential to recursively surpass the limitations associated with recurrence protocols based on Clifford operations. Notably, this enhancement persists even when operating on more than three copies, where only residual benefits have been observed with Clifford operations [41].
Consider the three-copies double selection protocol [39] illustrated in Fig. 1b. Following similar ideas to those used for the coherent extension of the two-copy protocol discussed in Sec. III.1, one can design a coherent version of the double selection protocol [39]. This entails integrating a control state that facilitates the construction of an effective $4\rightarrow 1$ protocol, see Fig. 6a. It is important to note that the dimension of the control copy is now greater than two, necessitating careful consideration when assessing resources and protocol efficiency.
The double selection strategy involves three identical copies, of which two are eventually measured in different bases. Consequently, there are two target copies and three different roles during the protocol. Initially, we consider a three-dimensional control qubit (qutrit) state enabling us to coherently permute the surviving copy role. Subsequently, we explore a six-dimensional control copy allowing coherent permutations among all possible role configurations.
@@ IV Further extensions
Notice that these choices are made for illustrative purposes. In general, one can also construct higher-dimensional control entangled copies using the same ensemble qubit EPR states, effectively building control states of dimension $2^{k}$, where $k$ represents the number of entangled copies involved.
Fig. 7 illustrates the performance of the aforementioned superposed double selection protocol, utilizing a higher dimensional Werner-like control copy with fidelity $F_{\rm c}=F^{log_{2}(D)}$, where $D$ is the qudit dimension, incorporating this dimensionality factor for a fair resource comparison. The controlled-SWAP (cSWAP) operation facilitates the coherent swapping of roles among the three states designating the surviving one in each branch. Following this, the double selection protocol continues as usual, with the control copy measured now in the generalized Fourier basis for $d=D$, and the desired outcomes post-selected.
@@
For comparison, we now include the triple selection protocol [41] (illustrated in Fig. 6b), a $4\rightarrow 1$ protocol that only yields marginal improvements compared to lower-copy Clifford approaches, see [41] and Fig. 7. However, our superposed strategy demonstrates a clear ability to surpass this marginality, showcasing its potential for significant enhancement.
In this case, we apply depolarization to Werner-like states after each round. Inspired by the analyses of the protocol discussed in Sec. III.1, one could expect further performance improvements by allowing for analogous variants and modifications.
V Conclusions
We have introduced a novel superposed entanglement purification and distillation strategy, leveraging coherent permutations of the roles of entangled states to enhance purification efficiency. Through comprehensive theoretical analyses and simulations, we have demonstrated the robustness and potential of our approach in overcoming limitations inherent in conventional entanglement purification techniques. Our findings indicate that the superposed strategy consistently achieves higher fidelities even in the presence of noise, surpassing traditional EPPs such as the double selection method.
Our findings suggest that the superposed approach, which allows building entanglement purification strategies with non-Clifford operations, can be hierarchically integrated with existing EPPs, consistently improving protocol performance.
Preliminary results suggest that the tools introduced in this work have the potential to enhance other variants of EPPs as well. Similar concepts could be also applied to different classes of entanglement purification strategies, such as hashing, where coherence tools could be utilized to accelerate convergence into the set of likely sequences, thereby improving protocol performance.
@@ Acknowledgments
Appendix A
We show here the performance of each entanglement purification protocol as a function of the evolution of each of the Pauli diagonal elements for different initial errors. In all protocols, the states remain Pauli diagonal during their execution.
One can observe how the $3\rightarrow 1$ protocols equal or outperform the basic Oxford one for all errors, while our superposed version clearly improves the double selection protocol for one of the errors and slightly outperforms it for initial Werner states.

### v2

@@ 3.2 Imperfect operations
The cSWAP operation, a non-Clifford gate also known as the Fredkin gate [59, 60], has been extensively explored in the literature, and various experimental proposals for its high-fidelity implementation already exist [61, 62, 63, 64, 65, 66, 67]. These proposals are based on alternative strategies that circumvent the limitations of decomposing the cSWAP gate into elementary gates. Alternatively, the superposed EPP scheme could be implemented in an interferometric-based way [52, 57], circumventing those problems and possibly enhancing the quality of the implementation of such a gate.
@@
Figs. 4a4b show the operational regime and yield of the protocols under operational amplitude damping noise, where the noise strength is assumed to be the same for the cNOTs ($q_{cnot}$) and cSWAPs operations ($q_{cSWAP}$). Notably, the superposed strategy, despite the noise from the cSWAP operations, tolerates higher levels of operational noise compared to the double selection protocol, both clearly surpassing more basic two-copy approaches. Additionally, the superposed EPP outperforms the double selection protocol in terms of fidelity evolution and overall performance.
Fig. 5a shows the operational regimes under operational local depolarizing noise and initial Werner states, for different values of the cSWAP noisy strength. We remark that the unique features of this gate could lead to alternative improved practical implementations [61, 62, 63, 64, 65, 66, 67, 52, 57].
Finally, we analyze in Fig. 5b the fidelity evolution for the different protocols under special operational noise considerations, motivated by realistic modeling of current quantum hardware [54]. Following such a model, operational noise is defined as a concatenation of local depolarizing, damping, and dephasing noise affecting each qubit after each gate, i.e., $\xi_{\rm damp}\circ\xi_{\rm damp}\circ\xi_{\rm deph}$. The same noise parameters are used for every gate, consistently showing an advantage of the superposed EPP.
Relevant and consistent performance improvements of the superposed EPP are found under different considerations of initial states and noisy operations.
@@
4 Further extensions
The tools presented in this work can offer a recursive method for constructing more efficient EPPs, which can be seamlessly integrated with existing entanglement purification strategies, by just implementing them in coherent superposition. This approach allows for the design of families of non-Clifford EPPs, where the controlled-swap operations serve as the non-Clifford extension of known protocols.
We illustrate and analyze various examples herein, emphasizing the advantage derived from our approach. Notably, this enhancement persists even when operating on more than three copies, a regime where previous studies [43] have shown only diminishing improvements in performance. Our results demonstrate that the superposed protocol continues to provide significant benefits beyond this point, offering a promising avenue for further optimization.
Consider again the double selection protocol [41] illustrated in Fig. 1b. Following similar ideas to those used for the coherent extension for the two-copy protocol discussed in Sec. 3.1, one can design a coherent version of the double selection protocol [41]. This can be done by integrating a control state that facilitates the construction of an effective $4\rightarrow 1$ protocol, see Fig. 6a. It is important to note that the dimension of the control copy is now greater than two, necessitating careful consideration when assessing resources and protocol efficiency.
The existing double selection strategy, Fig. 1b, involves three identical copies, of which two are eventually measured in different bases. Consequently, there are two target copies and three different roles during the protocol. For the superposed extension of such protocol, we first consider a three-dimensional control qubit (qutrit) state enabling us to coherently permute the surviving copy role. Subsequently, we explore a six-dimensional control copy allowing coherent permutations among all possible role configurations. Notice that these choices are made for illustrative purposes. In general, one can also construct higher-dimensional control entangled copies using the same ensemble qubit EPR states, effectively building control states of dimension $2^{k}$, where $k$ represents the number of entangled copies involved.
@@ 4 Further extensions
Fig. 7 shows the performance of the aforementioned superposed double selection protocol, utilizing a higher dimensional control copy initially prepared as a Werner state with fidelity $F_{\rm c}=F^{log_{2}(D)}$, where $D$ is the qudit dimension, incorporating this dimensionality factor for a fair resource comparison. The controlled-SWAP (cSWAP) operation facilitates the coherent swapping of roles among the three states designating the surviving one in each branch. Following this, the double selection protocol continues as usual, with the control copy measured now in the generalized Fourier basis for $d=D$, and the desired outcomes post-selected.
For comparison, we now include the triple selection protocol [43] (illustrated in Fig. 6b), a $4\rightarrow 1$ protocol (triple selection) that only yields marginal performance improvement compared to the double selection approach, see [43] and Fig. 7. Our superposed extension however demonstrates a clear ability to surpass this marginality, leading to further improvements that get even more relevant with higher dimension of the control register (i.e. larger number of superposed role exchanges).
@@
We note that in Fig. 7 analysis we apply depolarization to Werner-like states after each round of the protocol. However, inspired by the analyses of the protocol discussed in Sec. 3.1, one could expect further performance improvements by allowing for analogous variants and modifications as the ones discussed in Sec. 3.1.
5 Conclusions
In this work, we have introduced an alternative strategy for entanglement purification and distillation based on the coherent superposition of protocols. By leveraging coherent permutations of the roles of entangled states in existing protocols, we have shown how this approach consistently enhances other protocols performance. Through comprehensive theoretical analyses and numerical simulations for different initial states and under imperfect operations, we have demonstrated the robustness and potential of our approach. Our findings indicate that the superposed strategy achieves higher fidelities when using more copies, also in the presence of operational noise, surpassing traditional strategies such as the double and triple selection methods.
This approach also enables the design of alternative entanglement purification protocols that effectively utilize non-Clifford operations, expanding the landscape of entanglement distillation. Furthermore, it offers a hierarchical framework that can be integrated with existing purification strategies, with consistent performance enhancements.
Preliminary results suggest that the tools introduced in this work could benefit other variants of EPPs as well. For instance, similar concepts could be applied to hashing techniques, where coherence tools can improve convergence into the set of likely sequences, thereby boosting protocol efficiency. We hope this work can open new avenues in the designing of enhanced strategies for advancing in the field of entanglement purification.
@@ Acknowledgments
Appendix A. Density matrix evolution for different EPP protocols
We show here the performance of different entanglement purification protocol as a function of the evolution of each of the Pauli diagonal elements during protocol iterations. In all protocols, the states remain Pauli diagonal during their execution. We consider the Oxford protocol (Fig. 1 (a)), the double selection EPP (Fig. 1 (b)), and our superposed EPP (Fig. 2). Initial states considered vary for each row, including only Pauli $\sigma_{x},\sigma_{y},\sigma_{z}$ noise for the first three rows respectively, and initial Werner states in the last one.
One can observe how the $3\rightarrow 1$ protocols equal or outperform (i.e. converges faster to higher fidelities) the basic Oxford one for all initial errors, while our superposed version clearly improves the double selection protocol for one of the errors ($\sigma_{y}$) and slightly outperforms it for initial Werner states, while remain the same (in terms of fidelity evolution) for the other two Pauli errors ($\sigma_{x},\sigma_{z}$), although the density matrix evolution differ.